# Twisted crossed products of Banach algebras

### v1

@@ 1. Introduction
Since the late 1950s, the crossed product construction in C*-algebras has been an important tool both for constructing novel examples and for studying group representations. More generally, works on twisted actions of discrete groups on C*-algebras and on the C*-twisted crossed product can be traced back to the next decade in the classical papers [19, 28, 5]. Two main goals of this paper are: (1) to introduce a general definition of crossed products associated with twisted actions of locally compact groups on Banach algebras and (2) to recover certain $L^{p}$-operator algebra versions of well-known theorems for twisted crossed products of C*-algebras.
@@
Particular cases of twisted crossed products of Banach algebras are already present in [2] for discrete groups acting on locally compact Hausdorff spaces and also in [18, Section 3] where $F_{\lambda}^{p}(G,\sigma)$, the reduced twisted $L^{p}$-group algebra of a locally compact group $G$, is thoroughly studied and shown to be a full invariant for the pair $(G,\sigma)$ when $p\neq 2$. The general case of a full twisted crossed product of a general Banach algebra by a locally compact group has not yet been considered and involves more technical considerations concerning multiplier algebras. In this paper, we aim to fill this natural gap. To avoid technicalities, we opt to only consider Banach algebras that are nondegenerately represented on a Banach space and that have a contractive approximate identity (cai). Under this setting, for any twisted Banach algebra dynamical system $(G,A,\alpha,\sigma)$ (see Definition 3.1) we have obtained a universal crossed product $F_{\mathcal{R}}(G,A,\alpha,\sigma)$ associated to each family $\mathcal{R}$ of contractive covariant representations of the system. Specializing to the case when $A$ is an $L^{p}$-operator algebra and $\mathcal{R}$ consists of all contractive covariant representations on $L^{p}$-spaces, we give a $p$-analogue of the so-called Packer–Raeburn trick, by showing that, when $p\in(1,\infty)$, the full $L^{p}$-twisted crossed product of a $A$ by a locally compact group $G$ is “stably” isometrically isomorphic to an untwisted crossed product of the spatial tensor product of $A$ with the compact operators on $L^{p}(G)$.
@@
A brief history of the overall theory of crossed products for Banach algebras begins with the ordinary (i.e., untwisted) crossed products of Banach algebras by locally compact groups associated to different classes of covariant representations, which have been thoroughly studied in the series of (still unpublished [^1] Prof. Marcel de Jeu kindly informed us that these works were accepted on the condition that they be merged into a single memoir, a task that is still in progress.) papers [11, 9, 8]. A paper specializing only on full and reduced crossed products for Banach algebras admitting left bounded approximate identities was published shortly after, see [20]. At the same time, N.C. Phillips in [24] adapted the definitions in [11] to $L^{p}$-operator algebras, yielding definitions for $F^{p}(G,A,\alpha)$ and $F^{p}_{\mathrm{r}}(G,A,\alpha)$, the full and reduced crossed products of an $L^{p}$-operator algebra $A$ by a (second countable) locally compact group $G$. As a particular instance the $L^{p}$-group algebras $F^{p}(G)$ and $F^{p}_{\mathrm{r}}(G)$ are recovered when $G$ trivially acts on $\mathbb{C}$. It is worth mentioning that some C*-results do not always carry to the $L^{p}$-setting, such as Takai duality for $L^{p}$-crosssed products, which is shown to fail in general [26]. More recently, Phillips’ construction has also been generalized to étale groupoids [16] and to both $L^{p}$-twisted group algebras and reduced twisted groupoid $L^{p}$-operator algebras of an étale groupoid, see [18, 3, 1]. Other papers have appeared that deal with these algebras in various degrees of generality, such as [6] where the C*-core of the (reduced and untwisted) $L^{p}$-crossed products by discrete groups is studied, and [15] which deals with the computation of the topological stable rank and real rank for the (reduced) $L^{p}$-twisted crossed product of a discrete group $G$ of subexponential growth acting on the commutative algebra $C_{0}(X)$.
@@
In this work, we only work with nondegenerate Banach algebras (i.e., those that can be nondegenerately represented on a Banach space) that have a two sided cai. This is done in order to make sure that the multiplier algebra of such algebras is well behaved and still an $L^{p}$-operator algebra when the original Banach algebra is represented on an $L^{p}$-space. For the most part we let $G$ be any locally compact group, but starting in Section 5 we will require that $G$ is second countable, which allows us to use some results from [24]. In our definition of a twisted Banach algebra dynamical system $(G,A,\alpha,\sigma)$ (see Definition 3.1) we require the map $\alpha\colon G\to\mathrm{Aut}(A)$ to be strongly continuous and the twist $\sigma\colon G\times G\to\mathrm{Inv}_{1}(M(A))$ to be jointly strictly continuous. Thus, our assumptions on the twisted action are stronger than in [21, 22], where it is only required that both $\alpha$ and $\sigma$ are Borel maps. This simplifies many arguments, avoids any discussion of Polish groups related to Banach algebras, and it allows us to take results from [11, 9, 8, 24] in which also the stronger condition of continuity is assumed.
@@
Let $p\in(1,\infty)$, let $A\subseteq\mathcal{B}(L^{p}(\mu))$ be a separable nondegenerate $L^{p}$-operator algebra with a cai, let $\mathcal{K}_{G}^{p}=\mathcal{K}(L^{p}(G))$, and let $(G,A,\alpha,\sigma)$ be a twisted dynamical system. Then there is an action $\beta$ of $G$ on $\mathcal{K}_{G}^{p}\otimes_{p}A$ such that $\mathcal{K}_{G}^{p}\otimes_{p}F^{p}(G,A,\alpha,\sigma)$ is isometrically isomorphic to the untwisted crossed product $F^{p}(G,\mathcal{K}_{G}^{p}\otimes_{p}A,\beta)$.
@@ 2. Preliminaries
    AA is nondegenerately represented. That is, there is a Banach space $E$ and an isometric algebra homomorphism $\pi\colon A\to\mathcal{B}(E)$ such that
    $\pi(A)E\coloneqq\mathrm{span}\{\pi(a)\xi\colon a\in A,\xi\in E\}$
    is a dense subspace of $E$. We often will simply say that $A$ is nondegenerate.
Recall that Assumption (A.1) is enough to isometrically identify $A$ with a closed two-sided ideal of $M(A)$, the multiplier algebra of $A$ defined as the algebra of double centralizers for $A$ (see Section 3 in [4] for the general definition and properties of $M(A)$). Indeed, this is done via the map $\iota_{A}\colon A\to M(A)$ defined as
@@
where $L_{a}(b)\coloneqq ab$ and $R_{a}(b)\coloneqq ba$ for all $a,b\in A$ (see Lemma 3.3. in [4] for complete details).
It is also well known that Assumptions (A.1) and (A.2) above imply that $\pi$ extends uniquely to a contractive unital algebra homeomorphism
(2.2) $\widehat{\pi}\colon M(A)\to\mathcal{B}(E)$
satisfying $\widehat{\pi}\circ\iota_{A}=\pi$ (see Theorem 4.1 and Remark 4.2 in [17]). This guarantees an isometric isomorphism
@@
(see Corollary 3.6 in [4]). Therefore, after identifying $A$ with its isometric copy $\pi(A)$ in $\mathcal{B}(E)$, each $t\in M(A)$ naturally gives two maps $A\to A$, each one with norm $\|t\|$, defined by $a\mapsto ta\in A$ and $a\mapsto at\in A$. Thus, in this case $M(A)$ is isometrically identified with a subalgebra of $\mathcal{B}(A)$ via the map $t\mapsto(a\mapsto ta)$. Furthermore, if $u\in M(A)$ is invertible, we get a map $\mathrm{Ad}(u)\colon A\to A$ defined by
@@
Let $p\in[1,\infty)$. A Banach algebra $A$ is an $L^{p}$-operator algebra if there is a measure space $(X,\mathfrak{M},\mu)$ such that $A$ is representable on $L^{p}(\mu)$. That is, there is an isometric algebra homomorphism $A\to\mathcal{B}(L^{p}(\mu))$. If $\varphi$ is nondegenerate, we say $A$ is a nondegenerate $L^{p}$-operator algebra.
@@
Let $p\in[1,\infty)$, let $(\Omega,\mathfrak{M},\mu)$ be a measure space, and let $E$ be a Banach space $E$. The Lebesgue-Bochner $L^{p}$-space is
$L^{p}(\mu;E)\coloneqq L^{p}(\mu,\Omega;E)\coloneqq\{\xi\colon\Omega\to E\mbox{ measurable}\colon\textstyle{\int_{\Omega}}\|\xi(\omega)\|^{p}\ d\mu(\omega)<\infty\}.$
@@
For $p\in[1,\infty)$, there is a tensor product, called the spatial tensor product and denoted by $\otimes_{p}$. We refer the reader to Section 7 of [10] for complete details on this tensor product. We only describe below the properties we will need. If $(\Omega_{0},\mathfrak{M}_{0},\mu_{0})$ is a measure space and $E$ is a Banach space, then there is an isometric isomorphism
@@ 3. Twisted Banach Crossed Products
Very much in the sense of [5], we now define a twisted Banach $L^{1}$-algebra associated to a twisted action of a locally compact group $G$ on $A$. To do so, we let $\mathrm{Aut}(A)$ be the group of isometric automorphisms of $A$ (i.e., $\varphi\in\mathrm{Aut}(A)$ if $\varphi\colon A\to A$ is an isometric invertible algebra homomorphism) and $\mathrm{Inv}_{1}(M(A))$ be the group of invertible norm $1$ elements in $M(A)$.
@@
Let $(e_{\lambda})_{\lambda\in\Lambda}$ be a cai for $A$. We claim that $\big((1_{G},e_{\lambda})\big)_{\lambda\in\Lambda}$ is a strong cai for $\mathcal{A}$ in the sense of Definition 2.11 in Chapter VIII of [13]. That is, $\|(1_{G},e_{\lambda})\|_{\pi^{-1}(1_{G})}\leq 1$ and for every compact subset $Q\subseteq G\ltimes_{\alpha,\sigma}A$ and any $\varepsilon>0$ there is $\lambda_{0}\in\Lambda$ such that
@@
Hence, the claim follows at once from the fact that $(e_{\lambda})_{\lambda\in\Lambda}$ is a (strong) cai for $A$ and two standard compactness arguments. In fact, the details are exactly as the ones one will need to carry in the untwisted case (see 4.3 in Chapter VIII of [13] ), for the two displayed multiplications above are independent of the twist chosen.
Next, let $L^{1}(G\mid\mathcal{A})$ be the cross-sectional algebra of $\mathcal{A}$ as defined in Chapter VIII Section 5 of [13]. Since $\mathcal{A}$ has a strong cai, we apply Theorem 5.11 in Chapter VIII of [13] and get a (strong) cai for $L^{1}(G\mid\mathcal{A})$ that belongs to $C_{c}(G\mid\mathcal{A})$, the space of continuous sections $G\to G\ltimes_{\alpha,\sigma}A$ with compact support. Furthermore, by Remark 5.12 in Chapter VIII of [13], the net obtained is also a cai for $C_{c}(G\mid\mathcal{A})$ when equipped with the inductive limit topology (see II-14.3 in [12]). Finally, observe that the map $L^{1}(G,A,\alpha,\sigma)\to L^{1}(G\mid\mathcal{A})$, $f\mapsto\widetilde{f}$, where $\widetilde{f}\colon G\to G\ltimes_{\alpha,\sigma}A$ is given by
@@
Once again as in Remark 3.3, and to contrast with [21, Definition 2.3], we opt to follow [11, 24] and impose the stronger assumptions of continuity for $u$ rather that simply asking for it to be Borel.
@@
We are now ready to define twisted Banach crossed products, much in the spirit of Definitions 3.1 and 3.2 in [11].
@@
Finally, by Proposition 3.6, $(f_{\lambda,U})_{\lambda\in\Lambda,U\in\mathcal{U}}$ is a cai for $C_{c}(G,A,\alpha,\sigma)$ equipped with the inductive limit topology, so Part (4) in Remark 3.13 implies that $(\tau_{\mathcal{R}}(f_{\lambda,U}))_{\lambda\in\Lambda,U\in\mathcal{U}}$ is a $C_{\mathcal{R}}$-approximate identity for $F_{\mathcal{R}}(G,A,\alpha,\sigma)$, as wanted. ∎

### v2

@@ 1. Introduction
Since the late 1950s, the crossed product construction in C*-algebras has been an important tool both for to construct novel examples and to study group representations. More generally, works on twisted actions of discrete groups on C*-algebras and on the C*-twisted crossed product can be traced back to the next decade in the classical papers [19, 28, 5]. Two main goals of this paper are: (1) to introduce a general definition of crossed products associated with twisted actions of locally compact groups on Banach algebras and (2) to recover certain $L^{p}$-operator algebra versions of well-known theorems for twisted crossed products of C*-algebras.
@@
Particular cases of twisted crossed products of Banach algebras are already present in [2] for discrete groups acting on locally compact Hausdorff spaces and also in [18, Section 3] where $F_{\lambda}^{p}(G,\sigma)$, the reduced twisted $L^{p}$-group algebra of a locally compact group $G$, is thoroughly studied and shown to be a full invariant for the pair $(G,\sigma)$ when $p\neq 2$. The general case of a full twisted crossed product of a general Banach algebra by a locally compact group has not yet been considered and involves more technical considerations concerning multiplier algebras. In this paper, we aim to fill this natural gap. To avoid technicalities, we opt to only consider Banach algebras that are nondegenerately represented on a Banach space and that have a contractive approximate identity (cai). Under this setting, for any twisted Banach algebra dynamical system $(G,A,\alpha,\sigma)$ (see Definition 3.1) we have obtained a universal crossed product $F_{\mathcal{R}}(G,A,\alpha,\sigma)$ associated to each family $\mathcal{R}$ of contractive covariant representations of the system. Specializing to the case when $A$ is an $L^{p}$-operator algebra and $\mathcal{R}$ consists of all contractive covariant representations on $L^{p}$-spaces, we give a $p$-analogue of the so-called Packer–Raeburn trick, by showing that, when $p\in(1,\infty)$, the full $L^{p}$-twisted crossed product of $A$ by a locally compact group $G$ is “stably” isometrically isomorphic to an untwisted crossed product of the spatial tensor product of $A$ with the compact operators on $L^{p}(G)$.
@@
A brief history of the overall theory of crossed products for Banach algebras begins with the ordinary (i.e., untwisted) crossed products of Banach algebras by locally compact groups associated to different classes of covariant representations, which have been thoroughly studied in the series of (still unpublished [^1] Prof. Marcel de Jeu kindly informed us that these works were accepted on the condition that they be merged into a single memoir, a task that is still in progress.) papers [11, 9, 8]. A paper specializing only on full and reduced crossed products for Banach algebras with left bounded approximate identities was published shortly after, see [20]. At the same time, N.C. Phillips in [24] adapted the definitions in [11] to $L^{p}$-operator algebras, yielding definitions for $F^{p}(G,A,\alpha)$ and $F^{p}_{\mathrm{r}}(G,A,\alpha)$, the full and reduced crossed products of an $L^{p}$-operator algebra $A$ by a (second countable) locally compact group $G$. As a particular instance the $L^{p}$-group algebras $F^{p}(G)$ and $F^{p}_{\mathrm{r}}(G)$ are recovered when $G$ trivially acts on $\mathbb{C}$. It is worth mentioning that some C*-results do not always carry to the $L^{p}$-setting, such as Takai duality for $L^{p}$-crossed products, which is shown to fail in general [26]. More recently, Phillips’ construction has also been generalized to étale groupoids [16] and to both $L^{p}$-twisted group algebras and reduced twisted groupoid $L^{p}$-operator algebras of an étale groupoid, see [18, 3, 1]. Other papers have appeared that deal with these algebras in various degrees of generality, such as [6] where the C*-core of the (reduced and untwisted) $L^{p}$-crossed products by discrete groups is studied, and [15] which deals with the computation of the topological stable rank and real rank for the (reduced) $L^{p}$-twisted crossed product of a discrete group $G$ of subexponential growth acting on the commutative algebra $C_{0}(X)$.
@@
In this work, we only work with nondegenerate Banach algebras (i.e., those that can be nondegenerately represented on a Banach space) that have a two sided cai. This is done in order to make sure that the multiplier algebra of such algebras is well behaved and still an $L^{p}$-operator algebra when the original Banach algebra is represented on an $L^{p}$-space. For the most part we let $G$ be any locally compact group, but starting in Section 5 we will require $G$ to be second countable, which allows us to use some results from [24]. In our definition of a twisted Banach algebra dynamical system $(G,A,\alpha,\sigma)$ (see Definition 3.1) we require the map $\alpha\colon G\to\mathrm{Aut}(A)$ to be strongly continuous and the twist $\sigma\colon G\times G\to\mathrm{Inv}_{1}(M(A))$ to be jointly strictly continuous. Thus, our assumptions on the twisted action are stronger than in [21, 22], where it is only required that both $\alpha$ and $\sigma$ are Borel maps. This simplifies many arguments, avoids any discussion of Polish groups related to Banach algebras, and it allows us to take results from [11, 9, 8, 24] in which also the stronger condition of continuity is assumed.
@@
Let $p\in(1,\infty)$, let $A$ be a nondegenerate separably representable $L^{p}$-operator algebra with a cai, let $\mathcal{K}_{G}^{p}=\mathcal{K}(L^{p}(G))$, and let $(G,A,\alpha,\sigma)$ be a twisted dynamical system. Then there is an action $\beta$ of $G$ on $\mathcal{K}_{G}^{p}\otimes_{p}A$ such that $\mathcal{K}_{G}^{p}\otimes_{p}F^{p}(G,A,\alpha,\sigma)$ is isometrically isomorphic to the untwisted crossed product $F^{p}(G,\mathcal{K}_{G}^{p}\otimes_{p}A,\beta)$.
@@ 2. Preliminaries
Notation 2.3.
A representation of a Banach algebra $A$ on a Banach space $E$ will always mean a bounded algebra homomorphism $\pi\colon A\to\mathcal{B}(E)$. We say
  1. (1)
    $\pi$ is contractive when $\|\pi\|\leq 1$,
  2. (2)
    $\pi$ is isometric when $\|\pi(a)\|=\|a\|$ for all $a\in A$,
  3. (3)
    $\pi$ is spearable when $E$ is a separable Banach space,
  4. (4)
    $\pi$ is nondegenerate when $\pi(A)E\coloneqq\mathrm{span}\{\pi(a)\xi\colon a\in A,\xi\in E\}$ is a dense subspace of $E$,
  5. (5)
    $\pi$ is unital when $A$ is unital and $\pi(1_{A})=\mathrm{id}_{E}$.
We say $A$ is representable on the Banach space $E$ when there is an isometric representation of $A$ on $E$.
@@
    AA is nondegenerately representable. That is, there is a Banach space $E$ and an isometric nondegenerate representation of $A$ on $E$. We often will simply say that $A$ is nondegenerate.
Recall that Assumption (A.1) is enough to isometrically identify $A$ with a closed two-sided ideal of $M(A)$, the multiplier algebra of $A$ defined as the algebra of double centralizers for $A$ (see [4, Section 3] for the general definition and properties of $M(A)$). Indeed, this is done via the map $\iota_{A}\colon A\to M(A)$ defined as
@@
where $L_{a}(b)\coloneqq ab$ and $R_{a}(b)\coloneqq ba$ for all $a,b\in A$ (see [4, Lemma 3.3] for complete details).
It is also well known that Assumption (A.1) above implies that if $\pi$ is a nondegenerate representation of $A$ on $E$ (see Notation 2.3 (4)), then it extends uniquely to a unital nondegenerate representation of $M(A)$ on $A$,
(2.2) $\widehat{\pi}\colon M(A)\to\mathcal{B}(E),$
satisfying $\widehat{\pi}\circ\iota_{A}=\pi$ (see [17, Theorem 4.1, Remark 4.2] and [11, Theorem 6.1]). Furthermore, $\widehat{\pi}$ is contractive when $\pi$ is and $\widehat{\pi}$ is isometric when $\pi$ is. This guarantees, using Assumption (A.2), an isometric isomorphism
@@
as shown in [4, Corollary 3.6]. Therefore, after identifying $A$ with its isometric copy in $\mathcal{B}(E)$, each $t\in M(A)$ naturally gives two maps $A\to A$, each one with norm $\|t\|$, defined by $a\mapsto ta\in A$ and $a\mapsto at\in A$. Thus, in this case $M(A)$ is isometrically identified with a subalgebra of $\mathcal{B}(A)$ via the map $t\mapsto(a\mapsto ta)$. Furthermore, if $u\in M(A)$ is invertible, we get a map $\mathrm{Ad}(u)\colon A\to A$ defined by
@@
Let $p\in[1,\infty)$. A Banach algebra $A$ is an $L^{p}$-operator algebra if there is a measure space $(\Omega,\mathfrak{M},\mu)$ such that $A$ is represented on $L^{p}(\mu)$. That is, there is an isometric representation $\pi\colon A\to\mathcal{B}(L^{p}(\mu))$. If $\pi$ is nondegenerate, we say $A$ is a nondegenerate $L^{p}$-operator algebra.
@@
Let $p\in[1,\infty)$, let $(\Omega,\mathfrak{M},\mu)$ be a measure space, and let $E$ be a Banach space $E$. The Lebesgue-Bochner $L^{p}$-space of $(\Omega,\mu;E)$ is
$L^{p}(\mu;E)\coloneqq L^{p}(\Omega,\mu;E)\coloneqq\{\xi\colon\Omega\to E\mbox{ measurable}\colon\textstyle{\int_{\Omega}}\|\xi(\omega)\|^{p}\ d\mu(\omega)<\infty\}.$
@@
For $p\in[1,\infty)$, there is a tensor product, called the spatial tensor product and denoted by $\otimes_{p}$. We refer the reader to [10, Section 7] for complete details on this tensor product. We only describe below the properties we will need. If $(\Omega_{0},\mathfrak{M}_{0},\mu_{0})$ is a measure space and $E$ is a Banach space, then there is an isometric isomorphism
@@ 3. Twisted Banach Crossed Products
Very much in the sense of [5], we now define a twisted Banach $L^{1}$-algebra associated to a twisted action of a locally compact group $G$ on $A$. To do so, we let $\mathrm{Aut}(A)$ be the group of isometric automorphisms of $A$ (i.e., $\varphi\in\mathrm{Aut}(A)$ if $\varphi\colon A\to A$ is an isometric invertible algebra homomorphism) and $\mathrm{Inv}_{1}(M(A))$ be the group of invertible elements $u\in M(A)$ with $\|u\|=\|u^{-1}\|=1$.
@@
Let $(e_{\lambda})_{\lambda\in\Lambda}$ be a cai for $A$. We claim that $\big((1_{G},e_{\lambda})\big)_{\lambda\in\Lambda}$ is a strong cai for $\mathcal{A}$ in the sense of [13, VIII-Definition 2.11]. That is, $\|(1_{G},e_{\lambda})\|_{\pi^{-1}(1_{G})}\leq 1$ and for every compact subset $Q\subseteq G\ltimes_{\alpha,\sigma}A$ and any $\varepsilon>0$ there is $\lambda_{0}\in\Lambda$ such that
@@
Hence, the claim follows at once from the fact that $(e_{\lambda})_{\lambda\in\Lambda}$ is a (strong) cai for $A$ and two standard compactness arguments. In fact, the details are exactly as the ones one will need to carry in the untwisted case (see [13, VIII-4.3] ), for the two displayed multiplications above are independent of the twist chosen.
Next, let $L^{1}(G\mid\mathcal{A})$ be the cross-sectional algebra of $\mathcal{A}$ as defined in Chapter VIII Section 5 of [13]. Since $\mathcal{A}$ has a strong cai, we apply [13, VIII-Theorem 5.11] and get a (strong) cai for $L^{1}(G\mid\mathcal{A})$ that belongs to $C_{c}(G\mid\mathcal{A})$, the space of continuous sections $G\to G\ltimes_{\alpha,\sigma}A$ with compact support. Furthermore, by [13, VIII-Remark 5.12], the net obtained is also a cai for $C_{c}(G\mid\mathcal{A})$ when equipped with the inductive limit topology (see [12, II-14.3]). Finally, observe that the map $L^{1}(G,A,\alpha,\sigma)\to L^{1}(G\mid\mathcal{A})$, $f\mapsto\widetilde{f}$, where $\widetilde{f}\colon G\to G\ltimes_{\alpha,\sigma}A$ is given by
@@
Once again as in Remark 3.3, and to contrast with [21, Definition 2.3], we opt to follow [11, 24] and impose the stronger assumptions of continuity for $u$ rather than simply asking for it to be Borel.
@@
We are now ready to define twisted Banach crossed products, much in the spirit of [11, Definitions 3.1 and 3.2].
@@
Finally, by Proposition 3.6, $(f_{\lambda,U})_{\lambda\in\Lambda,U\in\mathcal{U}}$ is a cai for $C_{c}(G,A,\alpha,\sigma)$ equipped with the inductive limit topology, so Part (4) in Remark 3.13 implies that the net $(\tau_{\mathcal{R}}(f_{\lambda,U}))_{\lambda\in\Lambda,U\in\mathcal{U}}$ is a $C_{\mathcal{R}}$-approximate identity for $F_{\mathcal{R}}(G,A,\alpha,\sigma)$, as wanted. ∎